OpenStudy (anonymous):

a compound with the empirical formula CH2 has a molar mass of 70 g/mol. what is the molecular formula for this compound?

OpenStudy (anonymous):

????

OpenStudy (gtxmuqsit):

It's pentene C5H10

OpenStudy (gtxmuqsit):

@emanuelfarias

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

what is the empirical formula of compound composed of 3.25% hydrogen, 19.36% carbon, and 77.39% oxygen by mass? can u help me with this one too?

OpenStudy (anonymous):

@GTXMUQSIT

OpenStudy (gtxmuqsit):

@emanuelfarias give 5 minutes

OpenStudy (anonymous):

@GTXMUQSIT K i will

OpenStudy (gtxmuqsit):

It's CH2O3 @emanuelfarias

OpenStudy (gtxmuqsit):

@emanuelfarias do you want solution of it?

OpenStudy (anonymous):

@GTXMUQSIT because it give me HCO and it say insert the subscripts needed that is

OpenStudy (anonymous):

@GTXMUQSIT just the number u know

OpenStudy (gtxmuqsit):

use the molar weights and assume there is 100g 3.25/1.008 = 3.22 19.36/12 = 1.61 77.39/16 = 4.84 now take the smallest value and divide all others by it in this case 1.61 is smallest. 1.61/1.61 = 1 therefore C=1 3.22/1.61 = 2 therefore H=2 4.84/1.61 = 3 therefore H=3 @emanuelfarias I hope now you get it

OpenStudy (anonymous):

so then it would be H2C1O

OpenStudy (anonymous):

nope it wrong cuz i need the subscripts of H , C and O

OpenStudy (gtxmuqsit):

the last one I typed wrong it's not H it's O=3

OpenStudy (gtxmuqsit):

\[CH _{2}O _{3}\]

OpenStudy (gtxmuqsit):

CH2O3

OpenStudy (gtxmuqsit):